# Optical and acoustic plasmons in the layered material SrRuO I Abstract

### v1

@@ VI Perspectives
The present study has demonstrated that optical and acoustic plasmons can be investigated by T-EELS in the complete BZ in layered systems. Therefore, with the advent of higher energy resolution, T-EELS will be competitive at lower energies with RIXS, also taking into account that momentum-resolved T-EELS provides a direct probe of the dynamic susceptibility. Then, in "strange" metals, it will be possible to study low-energy and high-momentum charge excitations which were predicted in Ref. [32] to depend on correlation effects. In this way it will be possible to study the spacial dependence of the density-density fluctuations in "strange" metals.
@@ VII Methods
VII.1 Samples
SrRuO crystals were grown using the floating-zone method [33]. The superconducting transition temperature of the sample was 1.5 K.
Thin TEM lamellas of SrRuO with the normal pointing along were prepared by Focused Ion Beam (FIB) using a Thermofisher instrument. The target thickness of the lamellas was 80 nanometers. Low ion energy polishing was used as final step to thin Ga-ion damaged surface layers.
VII.2 EELS measurements
The momentum-resolved loss function was recorded at a FEI Titan TEM equipped with a Wien monochromator and a Gatan Tridiem imaging filter (GIF) at 80 kV acceleration voltage in a serial way (i.e., one EEL spectrum per fixed momentum). The energy resolution is around 120 meV (FWHM of zero loss peak). Using a camera length (i.e., effective distance between sample and detector) of 1.15 m. The GIF entrance aperture was used to select the different momentum values and covers a momentum range of 0.04 . Due to instabilities of the monochromator and the rather long collection times required at large momentum transfers, the recorded spectra are subject to substantial noise as well as mutual random fluctuations/offsets. In order to mitigate these effects, each spectrum was separately aligned with the help of the quasi-elastic peak. Subsequently the quasi-elastic peak was subtracted (the peak was obtained by reflecting the gain tail at zero loss energy), before analyzing the position of the plasmon peak.
VII.3 Fermi velocities and susceptibility
For the calculation of the Fermi velocity along the [100] direction and the bare particle susceptibility we used a TB band structure [28] based on a LDA band structure. We use a multi-band version of Eq. 2 and taking only intra-band excitations into account with the same matrix element, thus neglecting inter-band excitations. Mainly the and the bands contribute to and .
@@ Acknowledgements.
J.F. thanks P. Abbamonte, R. von Baltz, S.-L. Drechsler, and A. Greco for helpful discussions. This work is supported by a KAKENHI Grants-in-Aids for Scientific Research (Grants No. 18K04715, No. 21H01033, and No. 22K19093), and Core-to-Core Program (No. JPJSCCA20170002) from the Japan Society for the Promotion of Science (JSPS) and by a JST-Mirai Program (Grant No. JPMJMI18A3). JS received funding from the HORIZON EUROPE framework program for research and innovation under grant agreement n. 101094299. AL and MK acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)-project-id 461150024. BB received funding from the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-id 390858490).X Author contributions

### v2

@@ VI Perspectives
The present study has demonstrated that optical and acoustic plasmons can be investigated by T-EELS in the complete BZ in layered systems. Therefore, with the advent of higher energy resolution, T-EELS will be competitive at lower energies with RIXS, also taking into account that momentum-resolved T-EELS provides a direct probe of the dynamic susceptibility. It will be possible to study in more detail the different influence of correlation effects on optical and acoustic plasmons, caused by an energy dependent effective mass. The latter was predicted by a combined density functional/ dynamical mean-field theory (DFT + DMFT) calculation [51] and experimentally detected by optical spectroscopy [46]. Furthermore, in "strange" metals, it will be possible to study low-energy and high-momentum charge excitations which were predicted in Ref. [52] to depend on correlation effects. In this way it will be possible to evaluate the spatial dependence of the density-density fluctuations in "strange" metals.
@@ VII Methods
VII.1 Dielectric Response of the Layered Plasmon System
The dielectric function corresponding to the Fetter model of a system of coupled 2D layers reads [38]
This dielectric function is an approximation assuming a perturbation charge that is confined to the 2D layers supporting the plasmons, which is violated by the electron beam resulting in deviations to the experimental dielectric response. However, the intensities of the plasmon peaks can be derived from this dielectric function by calculating the loss probability (T-EELS signal) using Eq. 1 and integration along (see Fig. 6).
A version of the long wave wavelength sum rule for the dynamic susceptibility reads [50]
Thus, in the limit, the dynamic susceptibility is determined by the longitudinal plasmon.
VII.2 Samples
Sr2RuO4 crystals were grown using the floating-zone method [53]. The superconducting transition temperature of the sample was 1.5 K.
Thin TEM lamellas of Sr2RuO4 with the normal pointing along were prepared by Focused Ion Beam (FIB) using a Thermofisher instrument. The target thickness of the lamellas was 80 nanometers. Low ion energy polishing was used as final step to thin Ga-ion damaged surface layers.
VII.3 EELS measurements
The momentum-resolved loss function was recorded at a FEI Titan3 TEM equipped with a Wien monochromator and a Gatan Tridiem imaging filter (GIF) at 80 kV acceleration voltage in a serial way (i.e., one EEL spectrum per fixed momentum). The energy resolution is around 120 meV (FWHM of zero loss peak). Using a camera length (i.e., effective distance between sample and detector) of 1.15 m. The GIF entrance aperture was used to select the different momentum values and covers a momentum range of 0.04 (0.13 mrad semi collection angle). The acquisition times are presented in Table 1.
0.08 | 0.12 | 0.16 | 0.33 | 0.49 | 0.66 | ||
0 | 2 | 5 | 5 | 5 | 10 | 10 | |
0.1 | 5 | 5 | 5 | 5 | 10 | 10 | |
0.2 | 5 | 5 | 5 | 5 | 10 | 10 | |
0.3 | 10 | 10 | 10 | 10 | 10 | 10 | |
0.4 | 10 | 10 | 10 | 10 | 10 | 10 | . |
Due to instabilities of the monochromator and the rather long collection times required at large momentum transfers, the recorded spectra are subject to substantial noise as well as mutual random fluctuations/offsets. In order to mitigate these effects, each spectrum was separately aligned with the help of the quasi-elastic peak. After alignment of the zero loss position a superposition of the following three functions was fitted to the spectra. i) an asymmetric Pseudo-Voigt-profile ( with and ) reflecting the quasi elastic peak including ultra-low-loss excitations such as phonons, ii) a Drude-like function corresponding to the plasmon peaks and iii) a phenomenological background following a linear function. Finally, the spectral positions and half widths of the plasmons were derived from the fitted parameters and of the Drude function. The intensity (Fig. 6) corresponds to the area under the fitted Drude peak, which is obtained by integrating the latter over .
VII.4 Fermi velocities and susceptibility
For the calculation of the Fermi velocity along the [100] direction and the bare particle susceptibility we used a TB band structure [47] based on a LDA band structure. We use a multi-band version of Eq. 2 and taking only intra-band excitations with the same matrix element into account, thus neglecting inter-band excitations which may be caused by a strong spin-orbit coupling, which leads to a k-dependence of the orbital character of the bands [49]. Mainly the and the bands contribute to and .
VIII Data availability
All data supporting the findings are provided as figures in the article. Data files for all figures are available at https://opara.zih.tu-dresden.de//handle/123456789/763 and from the corresponding authors on request.
@@ Acknowledgements.
J.F. thanks P. Abbamonte, R. von Baltz, S.-L. Drechsler, and A. Greco for helpful discussions. This work is supported by a KAKENHI Grants-in-Aids for Scientific Research (Grants No. 18K04715, No. 21H01033, and No. 22K19093), and Core-to-Core Program (No. JPJSCCA20170002) from the Japan Society for the Promotion of Science (JSPS) and by a JST-Mirai Program (Grant No. JPMJMI18A3). J.S. received funding from the HORIZON EUROPE framework program for research and innovation under grant agreement n. 101094299. A.L. and M.K. acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)-project-id 461150024. B.B. received funding from the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-id 390858490).X Author contributions